A student is taking a 5 question T/F test. If the student chooses answers at random, what is the probability of getting all questions correct? I'd like to know the formula to solve.
1
Expert's answer
2012-07-06T08:30:24-0400
The probability of getting correct (or incorrect) answer in a T/F question is 1/2. So, the probability of getting all questions correct is

P = (1/2)^5 = 1/32.